# A Bayesian Optimization Approach to Machine Translation Reranking

### v1

@@ Abstract
Reranking a list of candidates from a machine translation system with an external scoring model and returning the highest-scoring candidate remains a simple and effective method for improving the overall output quality. Translation scoring models continue to grow in size, with the best models being comparable to generation models. Thus, reranking can add substantial computational cost to the translation pipeline. In this work, we pose reranking as a Bayesian optimization (BayesOpt) problem. By strategically selecting candidates to score based on a balance of exploration and exploitation, we show that it is possible to find top-scoring candidates when scoring only a fraction of the candidate list. For instance, our method achieves the same CometKiwi score using only 70 scoring evaluations compared a baseline system using 180. We present a multi-fidelity setting for BayesOpt, where the candidates are first scored with a cheaper but noisier proxy scoring model, which further improves the cost-performance tradeoff when using smaller but well-trained distilled proxy scorers.
footnotetext: Code available soon.
@@ 1 Introduction
The quality of machine translation (MT) generation models has surged in recent years due to innovations in neural network architecture (Stahlberg, 2020), training data augmentation (Sennrich et al., 2016; Edunov et al., 2018), and scale of training data (Ott et al., 2018). At the same time, translation quality evaluation models have made similar gains for similar reasons (Rei et al., 2020; Juraska et al., 2023; Sellam et al., 2020) in addition to the new abundance of quality estimation training data (Freitag et al., 2023b). Reranking a list of translation candidates with an external evaluation model, a technique with a long history in natural langauge processing (Collins and Koo, 2005; Charniak and Johnson, 2005), thus remains a relevant method for gaining additional performance from an MT system (Fernandes et al., 2022).
The term “reranking” is used to refer to any method which manipulates or replaces the scores assigned by a generation model. It typically refers to ranking an $n$-sized set of completed generations with an arbitrary scorer, although it can also be integrated into the generation process (Singhal et al., 2023). A wide array of scoring methods exist, including discriminative classifiers (Lee et al., 2021; Bhattacharyya et al., 2021), noisy channel decoding (Yee et al., 2019), and sampling-based Minimum Bayes Risk decoding (MBR; Eikema and Aziz, 2020). In each of these previous works, a list of candidate sequences is generated, rescored, and the highest-scoring candidate is returned.
In this work, we present a general-purpose reranking algorithm for MT which reduces the computational burden of scoring all $n$ candidates. Crucially, our aim is to support arbitrary scoring functions by using a combination of methods that have seen widespread adoption across many domains: Bayesian optimization (BayesOpt) and Gaussian processes (GP). BayesOpt (Shahriari et al., 2016) is a sequential black-box optimization method that uses the posterior mean and variance of unobserved data points to decide which points to evaluate next. GPs (Rasmussen and Williams, 2005) are flexible priors over functions with a tractable posterior that can effectively handle nonlinearity.
@@
We apply BayesOpt and GPs (BayesOpt+GP) to MT list reranking in a straightforward manner and show that it obtains close to the maximum achievable score with only a fraction of score evaluations. For example, when the maximal obtainable score across 200 randomly sampled candidates is 0.8216 CometKiwi, our method achieves 0.8210 with 70 score evaluations on average, while scoring 70 random candidates attains 0.8149, a difference of 0.0061 which is likely to be human-detectable according to (Kocmi et al., 2024b).
@@
Then, building upon previous works that use a faster but noisier proxy scoring function to prune the candidate list (Fernandes et al., 2022; Eikema and Aziz, 2022), we propose a multi-fidelity extension to BayesOpt which incorporates proxy scores to improve estimation. This is related to coarse-to-fine methods (Petrov, 2011) and model cascades (Kaynak and Alpaydin, 2000). Using this method, we find that smaller proxy scoring models distilled from the main model can further improve the performance of BayesOpt.
@@ 2.1 Translation generation and reranking
In a typical machine translation setting, a conditional language model is trained to model the probability of the next token $y_{t}$ given a source sentence $x$ and previous tokens: $p(y_{t}|x,y_{1},...,y_{t-1})$. These probabilities can be autoregressively combined to model a sequence probability $p(y|x)$. Usually, beam search is used to search for a $y$ which maximizes log probability combined with a length normalization objective (Wu et al., 2016).
In a basic list reranking setting, given $x$, the conditional language model is used to generate a candidate list $\mathcal{C}_{x}=[y_{1},...,y_{n}]$ with a decoding algorithm such as beam search or ancestral sampling. A scoring model $s(x,y_{i})$ is then applied to each $y_{i}\in\mathcal{C}_{x}$, and the best scoring sequence $\arg\max_{y_{i}\in\mathcal{C}_{x}}s(x,y_{i})$ is returned.
Reranking with high-quality evaluation metrics has been shown to be highly effective at improving translation output (Freitag et al., 2022), though it can skew results when the same metric is also used for evaluation (Kocmi et al., 2024a). Reranking performance improves as the number of candidates increases (Vernikos and Popescu-Belis, 2024) and when multiple scoring metrics are combined to form a stronger prediction (Fernandes et al., 2022). While such methods may be cost-prohibitive for large-scale deployment compared to beam search, they can be used in other parts of the MT development pipeline; high-quality predictions obtained from reranking are useful for knowledge distillation (Wang et al., 2024a) and self-training (Finkelstein et al., 2024).
Previous work on efficient reranking for MT is relatively limited. Fernandes et al. (2022) and Eikema and Aziz (2022) perform a two-stage reranking by first pruning with a faster and noisier scoring function to a fixed size before evaluating the target score. There has been recent interest in efficient approximations for MBR (Cheng and Vlachos, 2023; Deguchi et al., 2024; Trabelsi et al., 2024; Vamvas and Sennrich, 2024), but these methods do not generalize to other scoring functions.
@@ 2.2 Bayesian optimization withGaussian process prior
Bayesian optimization is a sequential algorithm for optimizing a black-box function $f$. In BayesOpt, $f$ is assumed to be drawn from a prior distribution over functions. The main loop of BayesOpt is as follows: given a set of (possibly noisy) observations of $f(a_{1}),...,f(a_{i})$, the prior distribution is updated to a posterior distribution with Bayes theorem. An acquisition function determines a query point $a_{i+1}$ at which to evaluate $f$ next. $f(a_{i+1})$ is evaluated and added to the set of observations. This repeats until a stopping criteria is reached. The principal design choices in BayesOpt are the prior distribution of $f$ and the acquisition function.
Our choice of prior is the Gaussian process, which assumes that any subset of points $f(a_{1}),...,f(a_{i})$ are drawn jointly from a multivariate Gaussian distribution $\mathcal{N}(\mu,\mathcal{K})$, where $\mathcal{K}$ is the covariance matrix defined by a kernel function such as the radial basis function kernel (RBF). RBFs define the covariance of two points $a$ and $a^{\prime}$ as:
@@
where $w$ is the bandwidth hyperparameter which determines scaling. The choice of kernel dictates prior assumptions about the shape of $f$; with RBF, points that are closer in Euclidean space have larger covariance. RBFs are a popular choice of kernel due their ability to model complex nonlinear functions.
@@
where $\mu$ is the unconditional mean of the distribution, $\sigma^{2}$ is a constant Gaussian noise on observations, and $I$ is the identity matrix.
@@ 3.1 MT reranking with Bayesian optimization
Our main algorithm is an adaptation of BayesOpt with GPs as described in Section 2.2 to the reranking setting. Each source sentence $x$ is treated as a standalone BayesOpt problem, meaning that no observations are shared across different $x$. Thus for brevity, we omit $x$ from notation when discussing BayesOpt for a particular instance.
Let $s$ be the scoring function, an MT quality estimator. Let $\mathcal{C}$ be a set of candidates, $\mathcal{C}_{\mathrm{obs}}\subseteq\mathcal{C}$ the subset of candidates for which we have observed $s(y)$, and $\bar{\mathcal{C}}_{\mathrm{obs}}$ be all other $y$ ($\bar{\mathcal{C}}_{\mathrm{obs}}=\mathcal{C}\setminus\mathcal{C}_{\mathrm{obs}}$). To perform reranking for a particular instance with candidates $\mathcal{C}$, we begin by scoring a small random subset of $\mathcal{C}$ with an MT quality estimator $s$. In one iteration in the algorithm loop, we normalize the observed scores to mean 0 and 1 variance at every step and assume a 0 unconditional mean. Then we compute the GP posterior of all $y\in\bar{\mathcal{C}}_{\mathrm{obs}}$ with Equation 2 and 3 given the scores of $\mathcal{C}_{\mathrm{obs}}$, which is then used to compute EI with Equation 5, assuming no observation noise. We evaluate the $k$ candidates in $\bar{\mathcal{C}}_{\mathrm{obs}}$ with the highest EI, removing them from $\bar{\mathcal{C}}_{\mathrm{obs}}$ and adding them to $\mathcal{C}_{\mathrm{obs}}$, and repeat the loop, terminating when a predefined budget of $n$ calls to $s$ is reached (or when all candidates have been evaluated, in the case that $|\mathcal{C}|\leq n$.). Finally, we choose $\arg\max_{y\in\mathcal{C}_{\mathrm{obs}}}s(y)$ as the prediction.
We now describe our choice of GP kernel. $y\in\mathcal{C}$ are strings, and we seek a representation that is fast to compute and to compare, since $|\mathcal{C}|$ representations are generated, and the computing the GP covariance matrix requires $|\mathcal{C}|^{2}$ comparisons. Our kernel is $\mathcal{K}_{\mathrm{MT}}(y_{i},y_{j})=\mathcal{K}_{\mathrm{RBF}}(\mathrm{emb}(y_{i}),\mathrm{emb}(y_{j}))$, where $\mathrm{emb}$ returns the mean-pooled token-level outputs of the final decoder layer when generating $y$, normalized to the unit norm after pooling. $\mathrm{emb}$ exploits meaning representations produced automatically during candidate list generation and thus requires negligible overhead. Also, the covariance matrix is fast to compute given the candidate list sizes and embedding dimensionality used in our experiments.
Inputs: scoring function $s$, proxy metric $s^{\prime}$, budget $n$ for evaluating $s$, covariance on dev set $\mathcal{K}_{\mathrm{score}}(s,s^{\prime})$,
          candidates $\mathcal{C}$, number of initial proxy scores $m$ and scores $n<m$, batch size $k$
Output: candidate $y\in\mathcal{C}_{\mathrm{obs}}$ with the highest score $s(y)$ among the observed candidates
1:$\mathcal{C}^{\prime}_{\text{obs}}\leftarrow\binom{\mathcal{C}}{m},\mathcal{C}_{\text{obs}}\leftarrow\binom{\mathcal{C}^{\prime}_{\text{obs}}}{n}$triangleright Score initial subsets
@@
[…]
5:while $|\mathcal{C}_{\mathrm{obs}}|<n$ and $|\mathcal{C}_{\text{obs}}|<|\mathcal{C}|$ do
@@
6:     $\hat{S}\leftarrow\text{Norm}(S_{\text{obs}}),\hat{S^{\prime}}\leftarrow\text{Norm}(S^{\prime}_{\text{obs}})$triangleright Normalize observed scores to 0-mean,1-variance
7:     $y_{\text{best}}\leftarrow\arg\max_{y\in\mathcal{C}_{\text{obs}}}\hat{S}(y)$
@@
8:     $\forall y\in\bar{\mathcal{C}}_{\mathrm{obs}}:\mu_{y},\sigma_{y}\leftarrow$ calculate posterior using $y,\mathcal{K}_{\mathrm{mult}},\hat{S},\hat{S^{\prime}}$triangleright GP posterior as in Equations 2 and 3
@@
9:     $\forall y\in\bar{\mathcal{C}}_{\mathrm{obs}}:\alpha_{y}\leftarrow\text{EI}(y_{\text{best}},\mu_{y},\sigma_{y})$triangleright Expected improvement as in Equation 5
@@
10:     $\mathcal{C}_{\text{top-}k}\leftarrow\operatorname{arg\,topk}_{y\in\bar{\mathcal{C}}_{\mathrm{obs}}}\alpha_{y}$triangleright Select k best candidates based on EI
@@
11:     $S_{\text{obs}}\leftarrow S_{\text{obs}}\cup\{s(y)|y\in\mathcal{C}_{\text{top-}k}\}$triangleright Compute scores for selected candidates
12:     $\mathcal{C}_{\text{obs}}\leftarrow\mathcal{C}_{\text{obs}}\cup\mathcal{C}_{\text{top-}k}$triangleright Update observed candidates
13:end while
14:return $\arg\max_{y\in\mathcal{C}_{\text{obs}}}s(y)$
  {algorithmB}
Algorithm for the proposed BayesOpt+GP+Proxy method. The proposed method selects a small subset of candidates, evaluates the true scoring function $s$ and the cheaper proxy metric $\hat{s}$ on initial subsets, and iteratively updates a Gaussian Process model using the the similarity between candidates’ embeddings and previously observed scores. The algorithm maximizes the Expected Improvement to select new candidates, updates the scores, and continues until the scoring budget $n$ is reached. If the proxy metric is not used and $\mathcal{K}_{\mathrm{mult}}=\mathcal{K}_{\mathrm{MT}}$, the algorithm reduces to BayesOpt+GP.

### v2

@@ Abstract
Reranking, or scoring a list of prediction candidates from a machine translation system with an external scoring model and returning the highest-scoring candidate, remains a simple and effective method for improving prediction quality. However, reranking with high quality scoring models can add substantial computational cost to the translation pipeline, which we address in this work by framing list reranking as a Bayesian optimization (BayesOpt) problem over the candidate list, where unknown scores are modeled with a Gaussian process. This algorithm scores candidates iteratively, choosing next candidates by balancing between exploration, choosing to score those that differ from candidates already scored, and exploitation, choosing to score those that resemble high-scoring candidates. This procedure finds high-scoring candidates while scoring only a fraction of the candidates list; given candidate lists of 200 random samples (before deduplication), our method achieves the same CometKiwi score using only 70 scoring evaluations on average compared to scoring a random subset of 180 candidates. We also propose multi-fidelity BayesOpt for list reranking, where scores obtained from a noisier but cheaper proxy scoring model are incorporated into the search process. We show that well-trained distilled proxy scorers can further improve the performance of BayesOpt.
footnotetext: Code repository: github.com/juliusc/bayesopt_reranking
@@ 1 Introduction
Reranking is a framework for prediction where probabilistic generator model produces a list of candidates, and a separate evaluator or scoring model produces scores for each of the candidates which are use to determine the final prediction. Reranking has a long history in natural language processing for sequential prediction problems such as dependency parsing (Collins and Koo, 2005; Charniak and Johnson, 2005) and language modeling problems such as summarization (Ravaut et al., 2022) and machine translation (MT; Fernandes et al., 2022).
The quality of models for automatic MT evaluation has surged in recent years due to innovations in neural network architecture (Rei et al., 2020; Juraska et al., 2023; Sellam et al., 2020) as well as the abundance of training data (Freitag et al., 2023b; Kocmi et al., 2024a). These evaluation models are often repurposed for reranking to further improve the performance of an MT system. For instance, in the WMT 2024 shared task (Kocmi et al., 2024a), 5 out of 19 systems, including the overall best submission (Rei et al., 2024) use reranking with Comet models (Rei et al., 2020) and/or minimum Bayes risk decoding (MBR; Eikema and Aziz, 2020), which can be interpreted as a form of reranking. Prior to the application of automatic evaluation metrics to reranking, other scoring methods have been proposed, including discriminatively trained classifiers (Lee et al., 2021; Bhattacharyya et al., 2021) and noisy channel decoding (Yee et al., 2019).
So, while LMs for MT generation for greatly improved in recent years, scoring models have seen a commensurate increase in quality (Zerva et al., 2022), and thus reranking remains relevant method for improving translation quality. However, the scoring models have also grown dramatically in size, increasing the computational requirements for reranking.
In this work, we address the computational cost of reranking by framing it as a search problem over the list of candidates. The goal of search in this setting is to find high-scoring candidates in a small number of steps, thereby avoiding the cost of scoring the full list. Our proposed algorithm uses Gaussian processes to model uncertainty about unseen scores and Bayesian optimization (BayesOpt; Shahriari et al., 2016) to choose which candidates to score next.
GPs are flexible priors over functions which are able to model the complex and nonlinear relationship between each candidate and its score. GPs make very few assumptions about the distribution and base their predictions are mostly on observed points, which enables them to easily adapt to different candidate lists across translation instances. BayesOpt is a sequential black-box optimization method that uses the posterior mean and variance of unobserved data points to decide which points to evaluate next.
@@
We apply BayesOpt and GPs (BayesOpt+GP) to MT list reranking in a straightforward manner and show that it obtains close to the maximum achievable score with only a fraction of score evaluations. For example, the maximal obtainable score across 200 randomly sampled candidates on our test set is 0.8216 CometKiwi; our method achieves 0.8210 with 70 score evaluations on average, while scoring 70 random candidates attains 0.8149, a difference of 0.0061 which is likely to be human-detectable according to (Kocmi et al., 2024b). We also propose a number of search-based baselines which outperform random selection, all of which are outperformed by BayesOpt+GP.
@@
Then, building upon previous works that use a faster but noisier proxy scoring function to prune the candidate list (Fernandes et al., 2022; Eikema and Aziz, 2022), we propose a multi-fidelity extension to BayesOpt which incorporates proxy scores to improve estimation. This is related in motivation to coarse-to-fine methods (Petrov, 2011) and model cascading (Chen et al., 2023), where the use of a faster proxy model reduces the use of the main model. In our multi-fidelity experiments, we find that smaller proxy scoring models distilled from the main model can assist BayesOpt+GP in finding high-scoring candidates earlier.
@@ 2.1 Translation generation and reranking
In a typical machine translation setting, a conditional language model (LM) is trained to model the probability of the next token $y_{t}$ given a source sentence $x$ and previous tokens: $p(y_{t}|x,y_{1},...,y_{t-1})$. These probabilities can be autoregressively combined to model a sequence probability $p(y|x)$. Usually, beam search is used to search for a $y$ which maximizes log probability combined with a length normalization objective (Wu et al., 2016).
In a basic list reranking setting, given $x$, the LM is used to generate a candidate list $\mathcal{C}_{x}=[y_{1},...,y_{n}]$ with a decoding algorithm such as beam search or ancestral sampling. A scoring function $s(x,y_{i})$ is then applied to each $y_{i}\in\mathcal{C}_{x}$, and the best scoring sequence $\arg\max_{y_{i}\in\mathcal{C}_{x}}s(x,y_{i})$ is returned. A common choice of scoring function is a quality estimation (QE) model which directly predicts a scalar value representing the quality.
Reranking with high-quality evaluation metrics has been shown to be highly effective at improving translation output (Freitag et al., 2022), though it can skew results when the same metric is also used for evaluation (Kocmi et al., 2024a). Reranking performance improves as the number of candidates increases (Vernikos and Popescu-Belis, 2024) and when multiple scoring metrics are combined to form a stronger prediction (Fernandes et al., 2022).
Reranking adds significant computational costs to prediction and may be prohibitive to use at test time, but it can be used to benefit LM training instead of test time prediction; high-quality predictions obtained from reranking can be used for knowledge distillation (Wang et al., 2024) and self-training (Finkelstein et al., 2024). Such methods can improve the performance of an MT system without additional costs during test time.
Previous work on efficient reranking for MT is relatively limited. Fernandes et al. (2022) and Eikema and Aziz (2022) perform a two-stage reranking by first pruning with a faster and noisier scoring function to a fixed size before evaluating the target score. There has been recent interest in efficient approximations for MBR (Cheng and Vlachos, 2023; Deguchi et al., 2024; Trabelsi et al., 2024; Vamvas and Sennrich, 2024), but these methods are not applicable to general scoring functions. (Singhal et al., 2023) propose to represent the candidate space compactly in a lattice over which a token-level reranker can efficiently score many candidates. In this work, we attempt to address a more general setting: the reranking of candidate lists with arbitrary black-box scoring functions.
@@ 2.2 Bayesian optimization withGaussian process prior
Bayesian optimization is a sequential algorithm for optimizing a black-box function $f$. $f$ is assumed to be drawn from a prior distribution over functions. The main loop of BayesOpt is as follows: given a set of (possibly noisy) observations of $f(a_{1}),...,f(a_{i})$, the prior distribution over $f$ is updated to a posterior distribution with Bayes theorem. An acquisition function determines a query point $a_{i+1}$ at which to evaluate $f$ next. $f(a_{i+1})$ is evaluated and added to the set of observations. This repeats until a stopping criteria is reached. The principal design choices in BayesOpt are the prior distribution of $f$ and the acquisition function.
A common choice of prior is the Gaussian process, which assumes that any subset of points $f(a_{1}),...,f(a_{i})$ are drawn jointly from a multivariate Gaussian distribution $\mathcal{N}(\mu,\mathcal{K})$, where $\mathcal{K}$ is the covariance matrix defined by a kernel function such as the radial basis function kernel (RBF). RBFs define the covariance of two points $a$ and $a^{\prime}$ as:
@@
where $w$ is the bandwidth hyperparameter which determines scaling. The choice of kernel dictates prior assumptions about the shape of $f$; with RBF, points that are closer in Euclidean space have larger covariance. RBFs are a popular choice of kernel due their ability to adapt to complex nonlinear functions.
@@
where $\mu$ is the unconditional mean of the distribution, $\sigma^{2}$ is a constant Gaussian noise on observations, $I$ is the identity matrix, and $\mathcal{K}$ here returns elementwise kernel values when given vector arguments.
@@ 3.1 MT reranking with Bayesian optimization
  Inputs: main metric $s$, proxy metric $s^{\prime}$, budget $n$ for evaluating $s$, hypotheses $\mathcal{C}$, number of initial main scores $\alpha$, number of initial proxy scores $\beta$, scoring budget $n$, batch size $k$, precomputed multi-fidelity kernel $\mathcal{K}_{mult}$.
Output: hypothesis with the highest observed score $\arg\max_{y\in\mathcal{C}_{\mathrm{obs}}}s(y)$.
1:$\mathcal{C}^{\prime}_{\text{obs}}\leftarrow\binom{\mathcal{C}}{\min(\beta,|\mathcal{C}|)},\mathcal{C}_{\text{obs}}\leftarrow\binom{\mathcal{C}^{\prime}_{\text{obs}}}{\min(\alpha,|\mathcal{C}|)}$triangleright Sample initial subsets
@@
4:while $|\mathcal{C}_{\mathrm{obs}}|<n$ and $|\mathcal{C}_{\text{obs}}|<|\mathcal{C}|$ do
5:     $\bar{\mathcal{C}}_{\mathrm{obs}}\leftarrow\mathcal{C}\setminus\mathcal{C}_{\mathrm{obs}}$triangleright Get complement of $\mathcal{C}_{\mathrm{obs}}$
@@
6:     $\hat{S}\leftarrow\text{Norm}(S_{\text{obs}}),\hat{S^{\prime}}\leftarrow\text{Norm}(S^{\prime}_{\text{obs}})$triangleright Normalize observed scores to 0 mean, 1 variance
7:     $y_{\text{best}}\leftarrow\arg\max_{y\in\mathcal{C}_{\text{obs}}}\hat{S}(y)$triangleright Get best observed point
@@
8:     $\forall y\in\bar{\mathcal{C}}_{\mathrm{obs}}:\mu_{y},\sigma_{y}\leftarrow$ calculate posterior using $y,\mathcal{K}_{\mathrm{mult}},\hat{S},\hat{S^{\prime}}$
9$\triangleright$ GP posterior as in Equations 2 and 3
@@
10:     $\forall y\in\bar{\mathcal{C}}_{\mathrm{obs}}:\gamma_{y}\leftarrow\text{EI}(y_{\text{best}},\mu_{y},\sigma_{y})$triangleright Expected improvement as in Equation 5
@@
11:     $\mathcal{C}_{\text{top-}k}\leftarrow\arg\mathrm{topk}_{y\in\bar{\mathcal{C}}_{\mathrm{obs}}}\gamma_{y}$triangleright Select $k$ best hypotheses based on EI
@@
12:     $S_{\text{obs}}\leftarrow S_{\text{obs}}\cup\{s(y)|y\in\mathcal{C}_{\text{top-}k}\}$triangleright Compute scores for selected hypotheses
13:     $\mathcal{C}_{\text{obs}}\leftarrow\mathcal{C}_{\text{obs}}\cup\mathcal{C}_{\text{top-}k}$triangleright Update observed hypotheses
14:end while
15:return $\arg\max_{y\in\mathcal{C}_{\text{obs}}}s(y)$
Algorithm 1 The BayesOpt+GP+P algorithm. BayesOpt+GP is a special case of this where $\beta=0$.
Our main algorithm is an adaptation of BayesOpt with GPs as described in Section 2.2 to the reranking setting. Each source sentence $x$ and its associated candidate list is treated as a standalone BayesOpt problem, meaning that no observations are shared across different $x$. Thus for brevity, we omit $x$ from notation when discussing BayesOpt for a particular instance.
Let $s$ be the scoring function, an MT quality estimator. Let $\mathcal{C}$ be a set of candidates, $\mathcal{C}_{\mathrm{obs}}\subseteq\mathcal{C}$ the subset of candidates for which we have observed $s(y)$, and $\bar{\mathcal{C}}_{\mathrm{obs}}$ be all other $y$ ($\bar{\mathcal{C}}_{\mathrm{obs}}=\mathcal{C}\setminus\mathcal{C}_{\mathrm{obs}}$). To perform reranking for an instance, we first generate candidates $\mathcal{C}$ and initialize the algorithm by scoring a random $\alpha$-sized subset of the list with $s$. In one iteration in the algorithm loop, we normalize the observed scores to mean 0 and 1 variance at every step and assume a 0 unconditional mean. Then we compute the GP posterior of all $y\in\bar{\mathcal{C}}_{\mathrm{obs}}$ with Equation 2 and 3 given the scores of $\mathcal{C}_{\mathrm{obs}}$, which is then used to compute EI with Equation 5, assuming no observation noise. We score the $k$ candidates in $\bar{\mathcal{C}}_{\mathrm{obs}}$ with the highest EI, adding them to $\mathcal{C}_{\mathrm{obs}}$ (as well as removing them from $\bar{\mathcal{C}}_{\mathrm{obs}}$), and repeat the loop, terminating when a predefined budget of $n$ calls to $s$ is reached (or when all candidates have been evaluated, in the case that $|\mathcal{C}|\leq n$.). Finally, we choose $\arg\max_{y\in\mathcal{C}_{\mathrm{obs}}}s(y)$ as the prediction.
We now describe our choice of GP kernel. $y\in\mathcal{C}$ are strings, and we seek a representation that is fast to compute and to compare, since $|\mathcal{C}|$ representations are generated, and the computing the GP covariance matrix requires $|\mathcal{C}|^{2}$ comparisons. Our kernel is $\mathcal{K}_{\mathrm{MT}}(y_{i},y_{j})=\mathcal{K}_{\mathrm{RBF}}(\mathrm{emb}(y_{i}),\mathrm{emb}(y_{j}))$, where $\mathrm{emb}$ returns the mean-pooled token-level outputs of the final decoder layer when generating $y$, normalized to the unit norm after pooling. $\mathrm{emb}$ uses meaning representations produced automatically during candidate list generation, so the additional cost to compute it is negligible. Also, the covariance matrix is fast to compute given the candidate list sizes and embedding dimensionality used in our experiments.